Circular No. 7449

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ct IN ANONYMOUS GALAXY S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000ct, obtained by A. Mahdavi on July 7.24 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type IIn supernova at an early epoch. The spectrum exhibits a very blue continuum, with strong Balmer emission lines as well as weak He I (rest 587.6 nm) emission. The Balmer lines consist of a narrow component (FWHM = 1000 km/s) on top of a broader base (FWHM = 3000 km/s). Cross-correlation of the host galaxy spectrum with an emission-line template yields a recession velocity of 8950 ± 30 km/s.

COMET C/1999 S4 (LINEAR) Recent binocular magnitude estimates: June 28.94 UT, 8.5 (A. Baran-

sky, Kiev, Ukraine, 12 × 80); 30.99, 8.3 (R. J. Bouma, Groningen, The Netherlands, 15 × 80); July 2.91, 7.7 (M. Reszelski, Szamotuly, Poland, 20 × 60); 5.06, 8.0 (S. Baroni, Milan, Italy, 40 × 80); 7.32, 8.0 (P. M. Raymundo, Salvador, Brazil, 11 × 80). T. Pauwels (Royal Observatory, Uccle, 0.85-m f/2.5 Schmidt + CCD) remarked on June 29.1 UT on tails of 14′ in p.a. 267^{o}, 8′ in p.a. 249^{o} and 4′ in p.a. 270^{o}. M. Jäger (Vienna, 0.25-m Schmidt, pg) remarked on July 7.0 on a type II tail extending 25′ in p.a. 267^{o} and type I tails extending 50′ in p.a. 275^{o} and 278^{o}.

MPEC 2000-N15 contains astrometric observations through July 6, new orbital elements and the ephemeris below. There is some indication that the comet’s motion is being influenced by nongravitational forces and that the comet could therefore be progressively shifted northward of this ephemeris, by perhaps up to 0′.1 by July 15 and up to 0′.6 by July 25.

2000TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 5 2^{h}49.^{m}63 +47^{o}06.^{′}4 0.846 0.869 54.^{o}7 72.^{o}7 8.0
7 2 59.16 +49 17.4 0.773 0.851 54.8  77.3 7.7
9 3 12.16 +51 50.7 0.701 0.835 54.5  82.3 7.4
11 3 30.66 +54 49.1 0.632 0.821 53.8  87.9 7.1
13 3 58.23 +58 11.6 0.565 0.807 52.4  93.9 6.8
15 4 41.26 +61 43.8 0.504 0.796 50.3 100.5 6.5
17 5 49.14 +64 34.5 0.450 0.786 47.5 107.5 6.2
19 7 24.57 +64 46.6 0.408 0.778 44.2 114.3 6.0
21 9 02.73 +60 17.7 0.381 0.772 41.2 119.8 5.8
23 10 15.01 +51 36.6 0.373 0.768 39.8 122.1 5.7
25 11 00.96 +41 08.6 0.385 0.765 40.6 120.3 5.8
2000 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden